# Leveraging Distribution Matching to Make Approximate Machine Unlearning Faster

### v1

@@ Abstract
Approximate machine unlearning (AMU) enables models to ‘forget’ specific training data through specialized fine-tuning on a retained dataset subset. However, processing this retained subset still dominates computational runtime, while reductions of epochs also remain a challenge. We propose two complementary methods to accelerate classification-oriented AMU. First, Blend, a novel distribution-matching dataset condensation (DC), merges visually similar images with shared blend-weights to significantly reduce the retained set size. It operates with minimal pre-processing overhead and is orders of magnitude faster than state-of-the-art DC methods. Second, our loss-centric method, Accelerated-AMU (A-AMU), augments the unlearning objective to quicken convergence. A-AMU achieves this by combining a steepened primary loss to expedite forgetting with a novel, differentiable regularizer that matches the loss distributions of forgotten and in-distribution unseen data. Our extensive experiments demonstrate that this dual approach of data and loss-centric optimization dramatically reduces end-to-end unlearning latency across both single and multi-round scenarios, all while preserving model utility and privacy. To our knowledge, this is the first work to systematically tackle unlearning efficiency by jointly designing a specialized dataset condensation technique with a dedicated accelerated loss function. Code is available at https://github.com/algebraicdianuj/DC_Unlearning.
@@ 1 Introduction
  • We benchmarked our approaches on various image classification datasets and models, achieving 84.61% and 54.45% faster end-to-end unlearning in single and multi-round scenarios, respectively, effectively filling latency gaps compared to current state-of-the-art AMU methods.
@@ 3.1 Preliminaries
Let the training dataset $\mathbf{D}$, consist of labeled images, denoted as $\{I_{p}\}_{p}$, comprising of $C$ classes and $N_{c}$ images per class, where $c\in\{1,2,\dots,C\}$. The training dataset be decomposed disjointly into retain $\mathbf{R}$ and forget $\mathbf{F}$ dataset. $\mathbf{D}=\{I_{p}\}_{p}$ was used to train a parameterized target network, and it achieved local minima parameters $\theta$ on average loss $\mathcal{L}(\theta;x,y)$ over $(x,y)\in\mathbf{R}\cup\mathbf{F}$. Let $\mathbf{D}$ and $\mathbf{\tilde{T}}$ be datasets drawn from the same underlying distribution $\mathcal{D}$, where $\mathbf{\tilde{T}}$ is the testing dataset, and is used in evaluation of training. Our objective is to unlearn forget set $\mathbf{F}\subset\{I_{p}\}_{p}$ from the pretrained model. Denote by $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ the empirical distribution of $\mathbf{R}$ and $\mathbf{F}$ respectively. The subset $\mathbf{T}\subseteq\mathbf{\tilde{T}}$ consists of elements whose class labels match those in $\mathbf{F}$, and $\mathcal{D}_{T}$ is its empirical distribution.
@@ 3.2 Blend Dataset Condensation
where $\omega_{ij}$ are scalar weights associated with the $j$-th image in sub-class $i$. These weights are treated as learnable parameters and are optimized to ensure that the feature representation of the blended image $\mathcal{I}_{i}$ closely matches the mean feature representation of its constituent images to form a condensed single image proxy of $\{I_{ij}\}_{j=1}^{n_{i}}$. We estimate $\omega_{i}$ in equation (2), via following distribution matching [29] loss:
@@ 3.3 Creating the Reduced Retain Dataset
Specifically, the inter-cluster indicator $\phi(j)$ and the intra-cluster indicator $\Phi(i)$ test whether the set of images $\{I_{ij}\}_{j=1}^{n_{i}}$ for each cluster $j$ intersects the forget set $\mathbf{F}$, thereby partitioning $R$ into $\mathbf{R}_{\text{free}}$ and $\mathbf{R}_{\text{residual}}$.
@@
6:return $\mathbf{R}_{\text{free}},\mathbf{R}_{\text{residual}}$
3.4 Accelerated Approximate Unlearning
$\mathbf{Q}^{(K)}(x)=\bigl{(}Q_{1}^{(K)}(x),\ldots,Q_{N}^{(K)}(x)\bigr{)}^{\!\top}\in(0,1)^{N}.$
3.4.1 Transforming losses of model to standard normal
Collect all losses of $\mathbf{F}$ and $\mathbf{T}$:
Applying the vector transformation in (4) to the loss vectors yields $\mathbf{Z}_{F}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{F})\bigr{)},\qquad\mathbf{Z}_{T}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{T})\bigr{)},$ where $\mathbf{Q}^{(K)}$ is the differentiable CDF estimator from Section 4.1 and $\Phi^{-1}$ is the probit of the standard normal distribution $\mathcal{N}(0,1)$. As $K\to\infty$ the mapping converges to the empirical probit, and this allows use to play with transformed losses in distribution matching manner. As emphasize previously, practice we keep $K$ sufficiently large to preserve differentiability.

### v2

@@ Abstract
Approximate machine unlearning (AMU) enables models to ‘forget’ specific training data through specialized fine-tuning on a retained dataset subset. However, processing this retained subset still dominates computational runtime, while reductions of epochs also remain a challenge. We propose two complementary methods to accelerate classification-oriented AMU. First, Blend, a novel distribution-matching dataset condensation (DC), merges visually similar images with shared blend-weights to significantly reduce the retained set size. It operates with minimal pre-processing overhead and is orders of magnitude faster than state-of-the-art DC methods. Second, our loss-centric method, Accelerated-AMU (A-AMU), augments the unlearning objective to quicken convergence. A-AMU achieves this by combining a steepened primary loss to expedite forgetting with a differentiable regularizer that matches the loss distributions of forgotten and in-distribution unseen data. Our extensive experiments demonstrate that this dual approach of data and loss-centric optimization dramatically reduces end-to-end unlearning latency across both single and multi-round scenarios, all while preserving model utility and privacy. To our knowledge, this is the first work to systematically tackle unlearning efficiency by jointly designing a specialized dataset condensation technique with a dedicated accelerated loss function. Code is available at https://github.com/algebraicdianuj/DC_Unlearning.
@@ 1 Introduction
  • Our benchmarks show 84.61% and 54.45% faster end-to-end unlearning in single and multi-round scenarios, respectively, on various image classification datasets and models, effectively filling latency gaps compared to current state-of-the-art AMU methods.
@@ 3.1 Preliminaries
Let the training dataset $\mathbf{D}$, consist of labeled images, denoted as $\{I_{p}\}_{p}$, comprising of $C$ classes and $N_{c}$ images per class, where $c\in\{1,2,\dots,C\}$. The training dataset is divided into retain $\mathbf{R}$ and forget $\mathbf{F}$ dataset. $\mathbf{D}=\{I_{p}\}_{p}$ was used to train a parameterized target network, and it achieved local minima parameters $\theta$ on average loss $\mathcal{L}(\theta;x,y)$ over $(x,y)\in\mathbf{R}\cup\mathbf{F}$. Let $\mathbf{D}$ and $\mathbf{\tilde{T}}$ be datasets drawn from the same underlying distribution $\mathcal{D}$, where $\mathbf{\tilde{T}}$ is the testing dataset, and is used in evaluation of training. Our objective is to unlearn forget set $\mathbf{F}\subset\{I_{p}\}_{p}$ from the pretrained model. Denote by $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ the empirical distribution of $\mathbf{R}$ and $\mathbf{F}$ respectively. The subset $\mathbf{T}\subseteq\mathbf{\tilde{T}}$ consists of elements whose class labels match those in $\mathbf{F}$, and $\mathcal{D}_{T}$ is its empirical distribution.
@@ 3.2 Blend Dataset Condensation
where $\omega_{ij}$ are scalar weights associated with the $j$-th image in sub-class $i$. These weights are treated as learnable parameters and are optimized to ensure that the feature representation of the blended image $\mathcal{I}_{i}$ closely matches the mean feature representation of its constituent images to form a condensed single image proxy of $\{I_{ij}\}_{j=1}^{n_{i}}$. We estimate $\omega_{i}$ in (2), via following distribution matching [29] loss:
@@ 3.3 Creating the Reduced Retain Dataset
Specifically, the inter-cluster indicator $\phi(j)$ and the intra-cluster indicator $\Phi(i)$ test whether the set of images $\{I_{ij}\}_{j=1}^{n_{i}}$ for each cluster $j$ intersects the forget set $\mathbf{F}$, thereby partitioning $\mathbf{R}$ into $\mathbf{R}_{\text{free}}$ and $\mathbf{R}_{\text{residual}}$.
@@
6:return $\mathbf{R}_{\text{free}},\mathbf{R}_{\text{residual}}$
3.4 Accelerated Approximate Unlearning
$\mathbf{Q}^{(K)}(x)=\bigl{(}Q_{1}^{(K)}(x),\ldots,Q_{N}^{(K)}(x)\bigr{)}^{\!\top}\in(0,1)^{N}$ (4)
3.4.1 Transforming losses of model to standard normal
Collect all losses of $\mathbf{F}$ and $\mathbf{T}$:
Applying the vector transformation in (5) to the loss vectors yields $\mathbf{Z}_{F}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{F})\bigr{)},\qquad\mathbf{Z}_{T}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{T})\bigr{)},$ where $\mathbf{Q}^{(K)}$ is the differentiable CDF estimator from (4) and $\Phi^{-1}$ is the probit of the standard normal distribution $\mathcal{N}(0,1)$. As $K\to\infty$ the mapping converges to the empirical probit, and this allows use to play with transformed losses in distribution matching manner. As emphasize previously, practice we keep $K$ sufficiently large to preserve differentiability.
3.4.2 Membership-Inference Attack (MIA) Loss
With the Gaussian kernel $\curlywedge(z,z^{\prime})=\exp\!\bigl{(}-\|z-z^{\prime}\|^{2}/2\bigr{)},$ the squared maximum mean discrepancy between the transformed loss distributions is
$\mathcal{L}_{\mathrm{MMD}}(\theta)$ $=\mathbb{E}_{z,z^{\prime}\sim\mathbf{Z}_{F}^{(K)}}\!\bigl{[}\curlywedge(z,z^{\prime})\bigr{]}+\mathbb{E}_{\tilde{z},\tilde{z}^{\prime}\sim\mathbf{Z}_{T}^{(K)}}\!\bigl{[}\curlywedge(\tilde{z},\tilde{z}^{\prime})\bigr{]}$ (6)
 $\quad-\,2\,\mathbb{E}_{z\sim\mathbf{Z}_{F}^{(K)},\,\tilde{z}\sim\mathbf{Z}_{T}^{(K}}\!\bigl{[}\curlywedge(z,\tilde{z})\bigr{]}.$
Minimising $\mathcal{L}_{\mathrm{MMD}}$ aligns the transformed loss distribution of the forget set with that of the class-matched test subset, thereby mitigating membership-inference leakage while remaining fully differentiable in $\theta$ and $K$.
3.4.3 Accelerated (A-) AMU Objective
We come to the second thesis of this paper, i.e. to accelerate unlearning at unlearning objective level. We propose following A-AMU minimizing objective, which proposes two main solutions over a general machine unlearning loss objective (mentioned in ‘Introduction’ section).
$\min_{\theta}\,\underbrace{\mathbb{E}_{(x,y)\sim\mathcal{D}_{R}}\!\bigl{[}\mathcal{L}(\theta;x,y)^{2}\bigr{]}}_{\textbf{Steeper Main Loss}}\;+\;\lambda\,\underbrace{\mathcal{L}_{\mathrm{MMD}}(\theta)}_{\textbf{Diff. MIA}}\;+\;\underbrace{\Gamma(\theta;\mathbf{R},\mathbf{F})}_{\textbf{Other Reg.}}$
where $\lambda\!>\!0$ is regularization parameter.
Solution # 1: Steeper main loss.
Replacing $\mathcal{L}$ by its square $\mathcal{L}^{2}$ multiplies the gradient magnitude by $2\,\mathcal{L}$. Consequently, gradient descent updates make larger progress early in training. Large errors on $\mathbf{R}$ are penalized disproportionately, while small errors become less important in course of optimization. So the model departs more rapidly from the decision boundaries, which it learned earlier during training schedule. In order to prevent overfitting on singular retain dataset samples for fixed learning rate, we take square of average loss per batch, instead of per sample squaring. This induces stability on choice of learning rate. Infact, in all of our experiments, we set a fixed learning rate for our approach on all models and datasets (including other hyperparameters), which is in strong contrast from other SOTA AMU methods that require hectic finetuning on each setting.
Solution # 2: Differentiable MIA term.
$\mathcal{L}_{\mathrm{MMD}}(\theta)$, from (6), essentially models something absolute in unlearning, i.e. for unlearned model, forget samples appears the same as test samples, which supersedes subjective modelling of forgetting regularizations in [15, 13, 6] dependent upon accuracy or parameter distribution hypothesis of unlearning. Minimizing this term over $\theta$ suppresses membership-inference signals on $\mathbf{F}$ while remaining fully differentiable in $\theta$. This interplays with making main loss steeper, as the parameter rapidly shift to overfit on retain set (or its reduced version), the differentiable MIA term allow quick restriction of parameters to only those trajectory, where MIA score is minimum.
Steeper main loss + Differentiable MIA term.
Combining the controlled instability from steeper main loss, around decision boundaries, with reduction of training trajectories towards absolute unlearning solutions, forms a strong acceleration of unlearning. As would seen later in our experiments, A-AMU objective minimization achieves competitive MIA suppression and retain accuracy after $\sim 1\,\text{epoch}$.
Additional Regularisation $\Gamma$.
The regularization term $\Gamma(\theta;\mathbf{R},\mathbf{F})$ serves as a characteristic placeholder for AMU unlearning ideas like weight sparsity [13], negative distillation [15], bad teacher distillation [6], etc.
4 Experiments
4.1 Experimental Setup
Methodology
We evaluate our methods on CIFAR-10, SVHN, and CINIC-10 datasets using ResNet-18 and ResNet-50 models. Our approach is benchmarked against six methods: Retraining, Catastrophic Forgetting (CF), Bad Distillation [6], $\mathbf{L_{1}}$-Sparsity [13], Pruning [13], and SCRUB [15]. For a fair comparison, implementations were adapted from their official repositories, and hyperparameters were rigorously fine-tuned for each model and dataset. [^2] Complete hyperparameter configurations are available in the ’hyperparameters’ folder of our code repository: https://github.com/algebraicdianuj/DC_Unlearning. We categorize these baselines into three tiers for analysis: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and our proposed accelerated (A-) methods.
Method w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC w/o DC w DC
A-CF $50.20\pm 2.65$ $52.27\pm 1.46$ $95.49\pm 4.84$ $97.06\pm 2.01$ $82.57\pm 3.42$ $84.83\pm 2.89$ $81.72\pm 5.24$ $84.03\pm 1.15$ $218.79\pm 0.34$ $133.32\pm 5.23$
Evaluation
Unlearning performance is assessed across five key metrics: Test Accuracy ($\uparrow$), Retain Accuracy ($\uparrow$), Forget Accuracy ($\downarrow$), Unlearning Time ($\downarrow$), and a MIA Score ($\downarrow$) for privacy. The MIA score is derived from a powerful white-box LiRA attack [3], providing a robust measure of privacy risk. For our Blend method, the $k$-means partitioning feature extractor is a lightweight CNN, 35×35$\times$35 × and 74×74$\times$74 × smaller than ResNet-18 and ResNet-50 respectively, ensuring its preprocessing overhead remains minimal. All experiments were conducted on a single NVIDIA RTX 3070 GPU (8GB VRAM), except for experiments in Sections 4.4 and 4.5, which were performed on a laptop RTX 3070 (8GB VRAM) due to computational resource availability during the extended experimental phase.
Figure 2: Qualitative and quantitative comparison of condensation methods on CIFAR-10/CNN. Top: class-wise synthetic prototypes generated by GM, DM, IDM, and our Blend method. Bottom: Performance metrics for each method over course of IPC = 2, 10.
Categorization:
We categorize baselines into three tiers: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and our accelerated (A-) which can contain acceleration of all naive and non-naive AMU methods.
4.2 The Efficiency and Efficacy of Blend Condensation
Our proposed dataset condensation method, Blend, is designed for both speed and efficacy, hence targeting unlearning. Its primary innovation is a reparameterization of the distribution matching objective that significantly reduces the number of trainable parameters compared to SOTA like Distribution Matching (DM), Gradient Matching (GM) and Improved Distribution Matching (IDM) [29, 30, 31]. For instance, on CIFAR-10 with an Images Per Class (IPC) of 10, Blend requires only $5\times 10^{4}$ parameters, whereas DM requires $30.72\times 10^{4}$.
We further shed light on this by conducting experiments on CNN and CIFAR-10. A key distinction in our setup is that Blend’s optimization process is limited to a synthetic data batch size of 1 in our implementation (as batched implementation is non-trivial). In contrast, competing methods like GM, DM, and IDM leverage large synthetic batch sizes (e.g., 256) for parallelized optimization. Despite being unable to use this common acceleration technique, Blend’s architectural efficiency leads to a dramatic performance gap. On CIFAR-10, Blend completes condensation in just 8-18 seconds (0.13–0.31 minutes). In stark contrast, GM, DM, and IDM require roughly 75–300 minutes, making our method over 1500x faster on average.
This dramatic speed-up does not compromise performance. As shown in Figure 2, Blend consistently yields higher test accuracy (e.g., 35.03% at IPC=10) compared to GM (33.01%), DM (32.2%), and IDM (33.1%). Furthermore, the synthetic prototypes generated by Blend are qualitatively superior—they are noticeably smoother and more class-consistent (e.g., clear airplane outlines) than the noisy, cluttered images from other methods. This demonstrates that our approach of grouping and blending feature-similar images is highly effective at producing coherent and informative exemplars.
@@ 4.2 The Efficiency and Efficacy of Blend Condensation
4.3 Single-Round Unlearning
[…]
Situational Impact of Blend
The effectiveness of Blend scales with the data reduction ratio. For random class removal on SVHN, shrinking the retain set to just 2.2% enabled the Pruning method to achieve a 96.3% total time reduction (from 647.4s to 23.7s) at the cost of only 3.9 percentage points in Test Accuracy, though the MIA score increased by 2.4. On CIFAR-10, a reduction to 20.0% of the retain set gave Retraining a 75.7% time reduction, but with a more substantial utility drop of 17.9 points in Test Accuracy and 24.5 points in Retain Accuracy. For 10% uniform forgetting on CIFAR-10 (60.2% retain), the gains were more modest: Retraining time was reduced by 37.8%, with a 6.2-point drop in Test Accuracy and a slight improvement in the MIA score (from 50.47 to 50.05). Thus, Blend offers dramatic speed-ups when data reduction is large, but this can introduce a trade-off with model utility.
Consistent Acceleration with A-AMU
Our A-AMU framework provides consistent algorithmic acceleration, independent of the unlearning scenario. On CIFAR-10 and without condensation, A-CF is 7.6x faster than naive CF for class removal (75.8s vs. 577.8s) and 7.4x faster for uniform forgetting (76.2s vs. 561.8s). This performance is achieved using a fixed set of hyperparameters for A-CF across all tasks.
Table 3: Unlearning on ResNet-18/SVHN (mean $\pm$ 95% CI) with $k$-means partitioning time $4.1379\pm 0.3221$ s, $\mathcal{F}$-sampling time $0.0537\pm 0.0089$ s